# Representation Meets Optimization: Training PINNs and PIKANs for Gray-Box Discovery in Systems Pharmacology

### v1

@@ 2.2 Multi-Layer Perceptron Representaion Model
The results so far for the forward problems regarding which representation is better, MLP or KAN, especially in the context of PINs are mixed and, in fact, problem-dependent [52].
@@ 2.3 Types of Error in Physics-Informed Networks
Physics-Informed Networks exhibit various types of errors, which can lead to reduced accuracy compared to traditional numerical methods. Figure 3 illustrates the sources of error during network training. The primary sources of these errors are as follows:
@@ 2.3.1 Approximation (Representation) Error
The first type of error in PINs is approximation (or representation) error, which for MLPs stems from the Universal Approximation Theorem. This theorem asserts that a feedforward neural network with sufficient neurons and a nonlinear activation function can approximate any continuous function $f:\mathbb{R}^{n}\to\mathbb{R}^{m}$ on a compact subset of $\mathbb{R}^{n}$. However, practical limitations such as network architecture, training time, and computational resources can lead to approximation errors in solving real-world physics-informed problems.
An alternative approach is the Kolmogorov-Arnold Network (KAN), inspired by the Kolmogorov-Arnold Representation Theorem (KART). KANs differ from MLPs by placing learnable activation functions on the edges (weights) instead of the nodes. This eliminates linear weight matrices and replaces each weight with a learnable 1D function, in this work modeled as a Chebyshev polynomial. The nodes in KANs simply sum incoming signals in each layer. Unlike MLPs, KART provides an exact representation but the approximation error is due to the parameterization of the univariate functions.
@@ 2.3.3 Optimization Error
In this context, let the differentiable loss function be denoted by $\ell:\mathbb{R}^{d}\to\mathbb{R}$, and its gradient as $\nabla\ell(\theta)$, where $\theta\in\mathbb{R}^{d}$ represents the model parameters. The goal of the optimization process is to find a parameter vector that minimizes the loss function, ideally reaching a local minimum.
@@
First-order optimization methods solve this problem by iteratively updating the parameter vector $\theta_{t}$ in a way that converges towards a local minimum. These methods rely on the evaluation of the loss function $\ell$ and its gradient $\nabla\ell$. The updates are guided by an update rule $\mathcal{M}$, which determines the next iterate $\theta_{t+1}$ based on the current history of iterates, gradients, and function values:
$\mathcal{H}_{t}=\{(\theta_{s},\nabla\ell(\theta_{s}),\ell(\theta_{s}))\}_{s=0}^{t}.$
@@
$\theta_{t+1}=\mathcal{M}(\mathcal{H}_{t},\phi_{t}),$
where $\phi_{t}$ is a set of hyperparameters, such as the learning rate, and the process starts from an initial value $\theta_{0}\in\mathbb{R}^{d}$.
Second-order optimization method
@@
where $\mathbf{p}_{k}$ is search direction pointing towards a region of lower function values on a high dimesnsion loss function and expressed as
@@
To construct a second-order optimizer, we need a combination of algorithms to compute both $\alpha_{k}$ and $\mathbf{p}_{k}$. In Equation 6, the inverse of the Hessian matrix $(\mathbf{H}_{k})^{-1}$ is required for updating the parameters. However, computing the Hessian matrix can be computationally expensive for high-dimensional problems. Moreover, away from the solution, the Hessian may not be positive definite and can become ill-conditioned. To balance computational efficiency and convergence performance, quasi-Newton methods [54] are employed. These methods iteratively approximate the Hessian matrix using only first-order derivative information, avoiding the need for explicit second-derivative computations. In this work, we use the BFGS update formula to compute $\mathbf{H}_{k}$, which is expressed as [55, 56]
@@
where $\mathbf{s}_{k}=\bm{\theta}_{k+1}-\bm{\theta}_{k}$ represents the change in the iterates, and $\mathbf{y}_{k}=\nabla\ell({\bm{\theta}_{k+1}})-\nabla\ell(\bm{\theta}_{k})$ represents the corresponding change in gradients and $\bm{v}_{k}=\sqrt{\bm{y}_{k}\cdot H_{k}\bm{y}_{k}}\left[\frac{\bm{s}_{k}}{\bm{y}_{k}\cdot\bm{s}_{k}}-\frac{H_{k}\bm{y}_{k}}{\bm{y}_{k}\cdot H_{k}\bm{y}_{k}}\right]$. $\tau_{k}$ and $\phi_{k}$ are are defined as the scaling and the updating parameters respectively. By choosing $\tau_{k}=\phi_{k}=1$, we recover the original BFGS algorithm [54]. The choice of $\tau_{k}$ and $\phi_{k}$ improves the conditioning of $\mathbf{H}_{k}$ and in this study we also compare the results against the Self-Scaled BFGS algorithm [56], for which $\tau_{k}$ and $\phi_{k}$ are chosen as
@@
In this study, we examine two line search approaches: 1) Backtracking, and 2) Trust region. The backtracking line-search strategy is based on the Armijo condition [57], which is formulated as follows,
@@
In the backtracking strategy, the step length is initially set to a reasonable value $\bar{\alpha}$ (not too small). If the condition is satisfied, the step size $\alpha_{k}$ is accepted (i.e., $\alpha_{k}=\bar{\alpha}$), and the algorithm proceeds to the next iteration. If the condition is not met, the step size is reduced by multiplying it by a factor $\rho<1$, and the process is repeated until the Armijo condition is satisfied. Additionally, the following condition is imposed along with (9),
@@
$&\min_{\mathbf{p}}m_{k}(\mathbf{p})=\nabla\ell_{k}^{\top}\mathbf{p}+\frac{1}{2}\mathbf{p}^{\top}\mathbf{H}_{k}\mathbf{p}\\ &\text{ subject to: }\\ &\|p\|\leq\Delta_{k}\end{aligned}$ (11)
@@
    Precision Settings: We investigate the impact of single-precision versus double-precision settings during training. Previous computational experiments [58] have shown that single-precision can achieve performance comparable to double-precision under certain conditions, particularly when the line search algorithm effectively identifies improvements. In this study, we aim to explore the circumstances under which double-precision settings outperform single-precision in solving inverse problems.
This analysis will provide insights into how various optimization techniques, along with precision settings, affect the training process and model performance when applied to PINNs and tanh-cPIKANs for gray-box discovery tasks.
@@ 3.1 Pharmacokinetics Model
Figure 4: Phamacokinetics model.
3.2 Phamacodynamics Model
The second model is a pharmacodynamics model used to describe the effect of a chemotherapy drug on cancer cell counts [59]. The time evolution of the cell count is governed by the following differential equation:
@@ 3.2 Phamacodynamics Model
Here, $r(D)$ is a dosage-dependent parameter that affects the time decay of the drug’s influence. These parameters are cell-line specific. Previous studies have fitted the functions $k_{d,A}(D)$ and $r(D)$ separately using nonlinear least squares methods, and the final model prediction for each dosage is a weighted combination of the results from both functions [59]. However, to simplify the computational model and avoid issues with identifiability, we assume that $k_{p}$ remains constant and that fitting both $k_{d}(D,t)$ and $\theta(D)$ simultaneously would lead to an unidentifiable solution. Therefore, we reduce the model to the simpler form as follows:
@@
For the simulations, we set the carrying capacity to $\theta(D)=1$, the intrinsic cell proliferation rate to $k_{p}=0.0345$, the maximum drug-induced death rate to $k_{d,B}=0.03$, and the drug efficacy decay rate to $r(D)=0.007$. These parameter values were chosen to qualitatively reflect typical tumor response dynamics under chemotherapeutic treatment, consistent with prior modeling studies [59]. The time-varying, nonlinear, and dose-dependent efficacy function $F_{D}(t)$ was designed to capture the dynamic impact of drug administration on cancer cell populations and serves as the ground truth to be identified through the inverse problem framework.
Figure 5: Phamacodynamics model.
@@ 4.1 Pharmacokinetics Model
Optimizers play an equally important role. First-order methods like Adam, RAdam, and SGD with momentum differ in how they handle gradient updates, adaptivity, and stability. Choosing an unsuitable optimizer can significantly degrade model performance. In this work, all experiments were implemented in JAX, and we employed the Optax and Optimistix [60] libraries for their efficient, modular, and JAX-compatible optimization algorithms.
For the pharmacokinetics model, both tanh-cPIKANs and PINNs used 2 hidden layers with 50 nodes/neurons per layer; this configuration resulted in approximately 10,300 trainable parameters for tanh-cPIKANs and 2,752 parameters for PINNs.
@@ 4.2.2 Performance of PINNs and tanh-cPIKANs with Single Precision
Overall, these findings emphasize that first-order optimizers, despite their slower progress, can sometimes reach better minima given enough iterations, while second-order methods excel in speed but may require double precision to realize their full potential.
@@ 4.3 Comparison against Self-Scaled BFGS
5 Summary
@@
In this work, we systematically investigated the optimization of two different physics-informed networks for gray-box discovery problems, focusing particularly on inverse problems in pharmacokinetics (PK) and pharmacodynamics (PD). Our aim was to answer key questions regarding model expressiveness, optimizer suitability, and training stability—especially in the context of ill-posed, sparse, and non-unique problems that arise frequently in PK/PD. To address these challenges, we introduced a slightly new architecture—tanh-cPIKANs—a variant of Chebyshev-based Kolmogorov–Arnold Networks (cPIKANs) with improved convergence and numerical stability. This design enhances gradient smoothness through the use of bounded activations (tanh), making it particularly effective for both first- and second-order optimization techniques (see also Appendix 2).
@@
We further showed that warm-up phases with first-order optimizers provide more robust training for second-order methods to reach optimal performance. Without warm-up, BFGS variants may converge prematurely to local minima due to poor initial curvature estimation—especially under single-precision arithmetic. This sensitivity underscores the importance of precision: double precision training significantly improves optimizer behavior, particularly for second-order methods that rely on gradient norms for convergence criteria.
@@
While our results provide valuable guidelines, several limitations remain. First, our study focused on moderately sized models and problems; the suitability of second-order methods for larger-scale PDE-constrained systems still requires careful memory and efficiency considerations. Second, although JAX enabled rapid prototyping and efficient auto-differentiation, its current limitations in mixed precision and full GPU memory utilization for large Hessian computations suggest further engineering optimizations are needed. Additionally, the performance of other KAN variants (e.g., rational or ReLU-based) was not investigated and is left for future work.
In conclusion, our findings demonstrate that optimizer choice, model architecture, and numerical precision are tightly coupled in gray-box discovery problems. The new tanh-cPIKANs represent a promising direction for scientific modeling tasks that demand expressiveness and smooth optimization landscapes. However, their full potential is only realized when paired with hybrid training strategies and sufficient numerical precision, reinforcing the importance of principled design choices in physics-informed learning pipelines.

### v2

@@ 2.2 Multi-Layer Perceptron Representation Model
The results so far for the forward problems regarding which representation is better, MLP or KAN, especially in the context of PINs are mixed and, in fact, problem-dependent [shukla2024comprehensive].
@@ 2.3 Types of Error in Physics-Informed Networks
PINs exhibit various types of errors, which can lead to reduced accuracy compared to traditional numerical methods. Figure 3 illustrates the sources of error during network training. The primary sources of these errors are as follows:
@@ 2.3.1 Approximation (Representation) Error
The first type of error in PINs is approximation (or representation) error, which for MLPs stems from the Universal Approximation Theorem. This theorem asserts that a feedforward neural network with sufficient neurons and a nonlinear activation function can approximate any continuous function $f:\mathbb{R}^{n}\to\mathbb{R}^{m}$ on a compact subset of $\mathbb{R}^{n}$. However, practical limitations such as network architecture, training time, and computational resources can lead to approximation errors in solving real-world physics-informed problems. KANs differ from MLPs by placing learnable activation functions on the edges (weights) instead of the nodes. This eliminates linear weight matrices and replaces each weight with a learnable 1D function, in this work modeled as a Chebyshev polynomial. The nodes in KANs simply sum incoming signals in each layer. Unlike MLPs, KANs provide an exact representation but the approximation error is due to the parameterization of the univariate functions.
@@ 2.3.3 Optimization Error
In this context, let the differentiable loss function be denoted by $\ell:\mathbb{R}^{d}\to\mathbb{R}$, and its gradient as $\nabla\ell(\bm{\theta})$, where $\bm{\theta}\in\mathbb{R}^{d}$ represents the model parameters. The goal of the optimization process is to find a parameter vector that minimizes the loss function, ideally reaching a local minimum.
@@
First-order optimization methods solve this problem by iteratively updating the parameter vector $\bm{\theta}_{k}$ in a way that converges towards a local minimum. These methods rely on the evaluation of the loss function $\ell$ and its gradient $\nabla\ell$. The updates are guided by an update rule $\mathcal{M}$, which determines the next iterate $\bm{\theta}_{k+1}$ based on the current history of iterates, gradients, and function values:
$\mathcal{H}_{k}=\{(\bm{\theta}_{s},\nabla\ell(\bm{\theta}_{s}),\ell(\bm{\theta}_{s}))\}_{s=0}^{k}.$
@@
$\bm{\theta}_{k+1}=\mathcal{M}(\mathcal{H}_{k},\phi_{k}),$
where $\phi_{k}$ is a set of hyperparameters, such as the learning rate, and the process starts from an initial value $\theta_{0}\in\mathbb{R}^{d}$.
Second-order (quasi-Newton) optimization method
@@
where $\mathbf{p}_{k}$ is search direction pointing towards a region of lower function values on a high dimension loss function and expressed as
@@
To construct a second-order optimizer, we need a combination of algorithms to compute both $\alpha_{k}$ and $\mathbf{p}_{k}$. In Equation 7, the inverse of the Hessian matrix $(\mathbf{H}_{k})^{-1}$ is required for updating the parameters. However, computing the Hessian matrix can be computationally expensive for high-dimensional problems. Moreover, away from the solution, the Hessian may not be positive definite and can become ill-conditioned. To balance computational efficiency and convergence performance, quasi-Newton methods [nocedal1999numerical] are employed. These methods iteratively approximate the Hessian matrix using only first-order derivative information, avoiding the need for explicit second-derivative computations. In this work, we use the BFGS update formula to compute $\mathbf{H}_{k}$, which is expressed as [AlBaali1993, AlBaali2005]
@@
where $\mathbf{s}_{k}=\bm{\theta}_{k+1}-\bm{\theta}_{k}$ represents the change in the iterates, and $\mathbf{y}_{k}=\nabla\ell({\bm{\theta}_{k+1}})-\nabla\ell(\bm{\theta}_{k})$ represents the corresponding change in gradients and $\bm{v}_{k}=\sqrt{\bm{y}_{k}\cdot H_{k}\bm{y}_{k}}\left[\frac{\bm{s}_{k}}{\bm{y}_{k}\cdot\bm{s}_{k}}-\frac{H_{k}\bm{y}_{k}}{\bm{y}_{k}\cdot H_{k}\bm{y}_{k}}\right]$. $\otimes$ represents the outer product. $\tau_{k}$ and $\phi_{k}$ are are defined as the scaling and the updating parameters respectively. By choosing $\tau_{k}=\phi_{k}=1$, we recover the original BFGS algorithm [nocedal1999numerical]. The choice of $\tau_{k}$ and $\phi_{k}$ improves the conditioning of $\mathbf{H}_{k}$ and in this study we also compare the results against the Self-Scaled BFGS algorithm [AlBaali2005], for which $\tau_{k}$ and $\phi_{k}$ are chosen as
@@
In this study, we examine two line search approaches: 1) Backtracking, and 2) Trust region. The backtracking line-search strategy is based on the Armijo condition [armijo1966minimization], which is formulated as follows,
@@
In the backtracking strategy, the step length is initially set to a reasonable value ( $\bar{\alpha}$; not too small). If the condition is satisfied, the step size $\alpha_{k}$ is accepted (i.e., $\alpha_{k}=\bar{\alpha}$), and the algorithm proceeds to the next iteration. If the condition is not met, the step size is reduced by multiplying it by a factor $\rho<1$, and the process is repeated until the Armijo condition is satisfied. Additionally, the following condition is imposed along with (9),
@@
$\begin{aligned} &\min_{\mathbf{p}}\nabla\ell_{k}^{\top}\mathbf{p}+\frac{1}{2}\mathbf{p}^{\top}\mathbf{H}_{k}\mathbf{p}\\ &\text{ subject to: }\\ &\|\bm{p}\|\leq\Delta_{k}\end{aligned}$ (11)
@@
    Precision Settings: We investigate the impact of single-precision versus double-precision settings during training. Previous computational experiments [double] have shown that single-precision can achieve performance comparable to double-precision under certain conditions, particularly when the line search algorithm effectively identifies improvements. In this study, we aim to explore the circumstances under which double-precision settings outperform single-precision in solving inverse problems.
Robustness to Observational Noise
To complement our optimization analysis, we evaluate the robustness of PINNs and tanh-cPIKANs when trained with noisy data. In real-world pharmacological settings, measurement noise is unavoidable; thus, it is critical to assess model stability under such perturbations.
@@ 3.1 Pharmacokinetics Model
Figure 4: Pharmacokinetics model.
3.2 Pharmacodynamics Model
The second model is a pharmacodynamics model used to describe the effect of a chemotherapy drug on cancer cell counts [chemodata]. The time evolution of the cell count is governed by the following differential equation:
@@ 3.2 Pharmacodynamics Model
Here, $r(D)$ is a dosage-dependent parameter that affects the time decay of the drug’s influence. These parameters are cell-line specific. Previous studies have fitted the functions $k_{d,A}(D)$ and $r(D)$ separately using nonlinear least squares methods, and the final model prediction for each dosage is a weighted combination of the results from both functions [chemodata]. However, to simplify the computational model and avoid issues with identifiability, we assume that $k_{p}$ remains constant and that fitting both $k_{d}(D,t)$ and $\theta(D)$ simultaneously would lead to an unidentifiable solution. Therefore, we reduce the model to the simpler form as follows:
@@
For the simulations, we set the carrying capacity to $\theta(D)=1$, the intrinsic cell proliferation rate to $k_{p}=0.0345$, the maximum drug-induced death rate to $k_{d,B}=0.03$, and the drug efficacy decay rate to $r(D)=0.007$. These parameter values were chosen to qualitatively reflect typical tumor response dynamics under chemotherapeutic treatment, consistent with prior modeling studies [chemodata]. The time-varying, nonlinear, and dose-dependent efficacy function $F_{D}(t)$ was designed to capture the dynamic impact of drug administration on cancer cell populations and serves as the ground truth to be identified through the inverse problem framework.
Figure 5: Pharmacodynamics model.
@@ 4.1 Pharmacokinetics Model
Optimizers play an equally important role. First-order methods like Adam, RAdam, and SGD with momentum differ in how they handle gradient updates, adaptivity, and stability. Choosing an unsuitable optimizer can significantly degrade model performance. In this work, all experiments were implemented in JAX, and we employed the Optax and Optimistix [optimistix] libraries for their efficient, modular, and JAX-compatible optimization algorithms.
For the pharmacokinetics model, both tanh-cPIKANs and PINNs used 2 hidden layers with 50 nodes/neurons per layer; this configuration resulted in approximately 10,300 trainable parameters for tanh-cPIKANs and 2,752 parameters for PINNs. In addition, in A.1 we provide results using architectures with a comparable number of parameters for both models. These experiments further demonstrate that the observed performance differences originate primarily from the representational and optimization characteristics of the tanh-cPIKAN architecture—namely, its smoother loss landscape and adaptive spectral representation—rather than differences in parameter capacity. By contrast, increasing the parameter count in PINNs (e.g., 50 neurons, 5 hidden layers, $\sim 10{,}400$ parameters) did not yield proportional gains in accuracy.
@@ 4.2.2 Performance of PINNs and tanh-cPIKANs with Single Precision
Overall, these findings emphasize that first-order optimizers, despite their slower progress, can sometimes reach better minima given enough iterations, while second-order methods excel in speed but may require double precision to realize their full potential. A comprehensive set of ablation studies examining the effects of architectural design and training strategies for cPIKAN, tanh-cPIKAN, and PINNs, as well as the impact of adaptive loss weighting for PINNs, is provided in Ablation Study for Pharmacokinetics Model — PINNs.
@@ 4.3 Comparison against Self-Scaled BFGS
4.4 Robustness of PINNs and tanh-cPIKANs under noisy observations
In this section, we investigate the robustness of PINNs and tanh-cPIKANs under varying levels of observational noise. Specifically, we examine how each model responds to 5% and 10% input noise. We focus on the impact of two second-order optimization methods, BFGS and SSBFGS, combined with two line search strategies: backtracking and trust-region. Our objective is to determine whether architectural choices and optimizer configurations influence the models’ ability to generalize from noisy data.
To simulate noisy measurements, we apply multiplicative uniform noise to the sampled outputs from the underlying ODE model. Let $X(t_{i})\in\{G(t_{i}),B(t_{i}),U(t_{i})\}$ denote the noise-free simulated values at time point $t_{i}$, and let $\eta(t_{i})\sim\mathcal{U}(-1,1)$ be independent uniform random variables. Then, the noisy observations $\tilde{X}(t_{i})$ are generated as:
$\tilde{X}(t_{i})=X(t_{i})\cdot\left(1+\epsilon\cdot\eta(t_{i})\right),$ (17)
where $\epsilon\in\{0.05,0.10\}$ denotes the noise level. This formulation perturbs each data point independently, introducing variability that mimics real-world measurement noise.
To quantify the impact of noise, we report both the absolute MAE values and their relative degradation compared to the noise-free baseline. The degradation factor is computed as the ratio between the MAE under noisy input and the MAE without noise for each method and optimizer configuration:
$\text{Degradation Factor}=\frac{\text{MAE}_{\text{noisy}}}{\text{MAE}_{\text{clean}}}.$ (18)
This metric helps assess how sensitively each model-optimizer pair responds to increasing levels of noise and highlights configurations that are more robust under uncertainty.
Model Optimizer MAE
noise 0% noise 5% noise 10%
tanh-cPIKANs BFGS_bck 5.92e-06 9.95e-05 (×16.8) 2.33e-04 (×39.4)
tanh-cPIKANs BFGS_trust 6.23e-06 9.51e-05 (×15.3) 1.19e-04 (×19.1)
tanh-cPIKANs SSBFGS_bck 5.92e-06 9.63e-05 (×16.3) 2.07e-04 (×35.0)
tanh-cPIKANs SSBFGS_trust 6.23e-06 6.29e-05 (×10.1) 1.39e-04 (×22.3)
PINNs BFGS_bck 2.26e-05 6.18e-04 (×27.3) 6.97e-04 (×30.8)
PINNs BFGS_trust 3.06e-05 4.73e-04 (×15.5) 6.87e-04 (×22.5)
PINNs SSBFGS_bck 1.42e-05 6.25e-04 (×44.0) 8.25e-04 (×58.1)
PINNs SSBFGS_trust 3.78e-05 7.84e-04 (×20.7) 9.12e-04 (×24.1)
Table 9: PK model: Evaluation of the robustness of PINNs and tanh-cPIKANs using different optimizers (BFGS and SSBFGS) in terms of MAE under noisy observation. All experiments were performed in double precision, using a cosine learning rate scheduler for the RAdam optimizer with an initial learning rate of 0.001. “BFGS-bck” refers to the BFGS optimizer with a backtracking line search method, while “BFGS-trust” refers to the BFGS optimizer using a trust-region line search method. Results are reported for two levels of input noise: 10% and 5%, applied independently to each method to assess robustness.
Table 9 presents the MAE results across clean and noisy settings. Overall, tanh-cPIKANs outperform PINNs in both accuracy and robustness to observational noise. At noise levels of $5\%$ and $10\%$, tanh-cPIKANs consistently achieve lower error rates and exhibit more stable relative degradation compared to their noise-free baseline. Notably, tanh-cPIKAN paired with SSBFGS_trust shows the smallest degradation, increasing only by 10.1× at $5\%$ noise and 22.3× at $10\%$ noise—representing the most stable performance among all configurations. Trust-region variants also improve robustness in other tanh-cPIKAN settings, with degradation remaining under 20× at $5\%$ noise across the board. In contrast, PINNs demonstrate significantly higher sensitivity to noise, especially with second-order optimizers. PINNs with SSBFGS_bck show the most severe error amplification, with 44.0× and 58.1× degradation at $5\%$ and $10\%$ noise, respectively. However, for some PINN configurations, the degradation from 5% to 10% noise is less drastic, likely because the error is already high at the 5% level—leaving less room for proportional increase. This saturation effect suggests that PINNs may hit a noise sensitivity ceiling earlier than tanh-cPIKANs. It is also worth noting that although tanh-cPIKAN with BFGS_bck maintains low absolute errors, it shows relatively higher degradation at $10\%$ noise (×39.4), likely due to the instability of backtracking-based line searches in high-noise regimes. This emphasizes that while model architecture plays a critical role, the choice of line search strategy can strongly influence robustness. These results confirm that robustness to observational noise is significantly enhanced by combining expressive architectures like tanh-cPIKANs with second-order optimizers equipped with conservative trust-region strategies. Such configurations offer a reliable pathway for building noise-resilient gray-box models in scientific applications.
5 Summary and Outlook
@@
In this work, we systematically investigated the optimization of two different physics-informed networks for gray-box discovery problems, focusing particularly on inverse problems in pharmacokinetics (PK) and pharmacodynamics (PD). While we used relatively simple PK/PD models and simulated data, our primary aim was to answer key questions regarding optimizer suitability and training stability— especially in the context of ill-posed, sparse, and non-unique problems that arise frequently in PK/PD. To address these challenges, we introduced a slightly new architecture—tanh-cPIKANs—a variant of Chebyshev-based Kolmogorov–Arnold Networks (cPIKANs) with improved convergence and numerical stability. This design enhances gradient smoothness through the use of bounded activations (tanh), making it particularly effective for both first- and second-order optimization techniques (see also A.3).
@@
We further showed that warm-up phases with first-order optimizers provide more robust training for second-order methods to reach optimal performance. Without warm-up, BFGS variants may converge prematurely to local minima due to poor initial curvature estimation—especially under single-precision arithmetic. This sensitivity underscores the importance of precision: double precision training significantly improves optimizer behavior, particularly for second-order methods that rely on gradient norms for convergence criteria. Although our study focuses on low-dimensional ODE-based PK/PD systems, the utilization of second-order and quasi-Newton methods can also be effective in large-scale settings such as PDEs, stiff dynamics, or high-dimensional parameter spaces. The scalability of second-order optimizers (those used in this study) primarily depends on the number of trainable parameters in the neural network. The dimensionality of the underlying PDE does not directly affect scalability because the PDE inputs can be sampled stochastically during each training iteration, keeping the effective computational load independent of the PDE dimension. Furthermore, even when the neural network contains a large number of parameters, the proposed optimizers can remain practical. This is because the optimizer does not require updating every parameter explicitly at each step; instead, it relies on gradient information and low-rank approximations to the Hessian. In addition, the Hessian (or its approximation) can be constructed using gradients computed for a randomly selected subset of parameters, which further reduces computational costs while still capturing sufficient curvature information to guide optimization effectively. As a result, the proposed optimizers can handle relatively high-dimensional parameter spaces more efficiently than full Newton-type methods while still retaining key second-order advantages. A recent complementary study [elham] shows that self-scaled quasi-Newton methods (e.g., SSBFGS, SSBroyden) remain highly effective for challenging PDE problems, but also identifies practical limitations related to memory growth and line-search sensitivity. We therefore note that while our optimization insights extend conceptually to more complex systems, their efficiency in large-scale regimes depends on memory-aware implementations and problem-specific numerical considerations.
@@
While our results provide valuable guidelines, several limitations remain. First, our study focused on moderately sized models and problems; the suitability of second-order or quasi-Newton methods for larger-scale PDE-constrained systems still requires careful memory and efficiency considerations. Second, although JAX enabled rapid prototyping and efficient automatic-differentiation, its current limitations in mixed precision and full GPU memory utilization for large Hessian computations suggest further engineering optimizations are needed. Additionally, the performance of other KAN variants (e.g., rational or ReLU-based) was not investigated and is left for future work.
In conclusion, our findings demonstrate that optimizer choice, model architecture, and numerical precision are tightly coupled in gray-box discovery problems. The new tanh-cPIKANs represent a promising direction for scientific modeling tasks that demand expressiveness and smooth optimization landscapes. Importantly, our robustness study further reveals that the interplay between architecture and optimization strategy also governs the model’s ability to handle noisy data—an essential consideration for real-world applications. The full potential of tanh-cPIKANs is realized when paired with hybrid training strategies, robust second-order optimizers, and sufficient numerical precision, reinforcing the importance of principled design choices in physics-informed learning pipelines.
Code and Data Availability
All codes, trained models, and synthetic PK/PD datasets used in this study are publicly available at https://github.com/NazAhmadi/PINNs_PIKANs_Graybox.